OpenStudy (anonymous):

the san paulo community poool can be can be filled in 12hr if water enters through a pipe alone or in 30hr if water enters through a hose alone. If water is entering both ways how long will it take to fill the poool?

OpenStudy (saifoo.khan):

POOL*

OpenStudy (anonymous):

23 hrs

OpenStudy (anonymous):

i got 4.8

OpenStudy (anonymous):

but i'm not very certain about that

OpenStudy (anonymous):

IDK

OpenStudy (anonymous):

I'm not certain about this but here we go. lets say you have 2 pools. one fills up to a volume v with a hose in 30 hours and the other of the same volume v with a pipe in 12 hour. If you fill the 2 pools (2v) sequentially (one after the other) 12+30=2v it will take you 42 hours to finish. 42=2v so How much would it have taken you if you combined them? 42/21=v

OpenStudy (anonymous):

The last line should read 42/2=v 21